# CLARA: A Modular Framework for Unsupervised Transit Detection Using TESS Light Curves

### v1

@@ 3.5 Controlled Parameter Space Exploration
We selected URF-4 for comprehensive parameter optimization based on its superior recall performance (36.6%) during our initial benchmarking evaluation (URF 1-4) and its inherent controllability through physically motivated synthetic generation. URF-4’s framework enables systematic exploration of four key synthetic parameters (and more not explored ones):
@@
Our analysis employs a systematic parameter sweep methodology, generating synthetic training sets for a URF4 sub-variant model by varying individual parameters while holding others constant at baseline values. Each parameter combination from the cartesian product space is used to train (3.2) independent URF models against a fixed real light curve dataset, resulting in $3\times 2\times 2\times 3=36$ sub-variant models.
@@ 3.6.2 Controlled Subset for Model Comparison
  • Trend Analysis Step: The performance of the 36 models on this small subset helps us analyze trends in their behavior in relation to the parameters of synthetic set generation and quantify those trends using the performance metrics in 3.6.4 and then test the generalization of those trends on independent sectors using cross sector evaluation of the models with more (and better, in terms of representation of quantification of model performance) performance metrics described in Table 3
@@ 3.6.3 TOI Importance Weighting
where SNR represents the signal-to-noise ratio of the planet detection, Depth (ppm) is the transit depth in parts per million, and Duration (days) is the transit duration in days. This heuristic emphasizes TOIs that are deeper, longer-duration, and higher signal-to-noise, making them both more detectable and more reliable for anomaly detection evaluation.
For fair comparison across sectors, TOI importance scores are normalized to the range $[0,1]$:
$\text{Normalized Importance}=\frac{\text{Importance Score}-\text{Score}_{\text{min}}}{\text{Score}_{\text{max}}-\text{Score}_{\text{min}}}$ (3)
@@ 3.6.4 Performance Metrics
3.7 Correlating Synthetic Parameters to Performance Metrics
3.7.1 Results (1)
Here, we present the systematic trends we observed for these 36 urf-4 sub-variant models. Figures 4, and 5 present the correlation between individual synthetic parameters and the performance metrics TOI Recall, Anomaly Rate and TOI Importance AUC (A.4).
Figure 4: ncurves, duration_hr (days A.3), noise (ppm) and cadence (minutes) clockwise from top left against TOI Recall (blue) and Anomaly Rate (red)
Figure 5: Synthetic features (ncurves, duration_hr(in days A.3), noise (ppm), cadence (minutes) - clockwise from top left) v.s TOI Importance AUC
To quantify the relative influence of synthetic training parameters on URF-4 performance, we employed Random Forest regression to model the relationship between our four key synthetic parameters (n_transits, duration_hr A.3, cadence_min, noise_ppm) and three primary evaluation metrics (TOI Recall, Anomaly Rate and TOI Importance AUC). This approach provides interpretable feature importance scores that reveal which synthetic design choices most strongly influence model behavior.
The Random Forest regression analysis, used to quantify how synthetic training parameters influenced URF-4 performance (Figure 6, 7), directly informed the finer-grained experiments presented in Section 4.1. In particular, noise (ppm) emerged as the most influential parameter affecting TOI Importance AUC and transit duration the most influential for anomaly rate and TOI Recall behavior. Motivated by this, we fixed other parameters and systematically varied noise and duration to study their specific impact on multiple performance metrics using a revised, multi-component scoring formulation 4.1 for $\alpha$ model selection for experiments in Section 4. This highlights how feature importance insights can guide targeted parameter space explorations and model selection strategies.
Figure 6: Feature importance comparison for URF-4 synthetic training parameters. Left panel shows importance scores for TOI Recall prediction, right panel for Anomaly Rate prediction. Transit duration (days A.3) emerges as the dominant factor for both metrics, followed by transit multiplicity. Random Forest regressors trained on 36-variant parameter sweep results with 42-seed random state for reproducibility.
Figure 8 presents the performance trajectories of the top 10 URF-4 variants ranked by TOI Recall AUC across varying anomaly threshold percentiles. It shows the top URF-4 optimal models achieving  40% recall in the top 25% of anomaly scores. However, while individual feature importance rankings guide parameter prioritization for focused optimization (4), the combined metric analysis 3.8 becomes essential for interpreting current results since different parameters optimize different performance aspects. The divergent trajectories between TOI Recall AUC and TOI Importance AUC in Figure 8 highlight the need for weighted combination scoring to balance detection completeness against astrophysical relevance in variant selection.
@@ 3.8 Combined Metric Analysis
$\text{S}_{\text{combined}}=\alpha\cdot\text{Recall AUC}_{\text{TOI}}+(1-\alpha)\cdot\text{Importance AUC}_{\text{TOI}}$ (5)
The weighting parameter $\alpha$ enables systematic control over the scientific trade-off between casting a wider net (high $\alpha$, emphasizing comprehensive TOI recall) versus casting a more accurate net (low $\alpha$, prioritizing astrophysically significant (higher values of importance metrics) targets). This parameterization allows researchers to tailor anomaly detection behavior to their specific discovery goals: survey completeness studies favor high $\alpha$ values to maximize detection rates, while targeted follow-up campaigns benefit from low $\alpha$ values that concentrate observational resources on the most promising candidates.
@@
    Convexity: Since $\alpha\in[0,1]$ $S_{\text{combined}}$ lies within the convex hull of $R$ and $I$. This guarantees that the combined score is always bounded between the two metrics and cannot extrapolate beyond them i.e. $S_{\text{combined}}\in[\min(R,I),\max(R,I)]$.
@@
In the second part of this research 4, we use the intersection of high alpha and low alpha models and use an improved anomaly scoring method - Normalized Weighted Room Sum Score (N-WRSS) - for quantifying anomaly candidates.
@@
To validate the stability and generalization of our optimal URF-4 configuration for $\alpha$=0.5, we evaluate the balanced model across multiple random subsets of unseen data compiled with stratified sampling from the same sector (sector 2). In this context, by "generalization" we refer to how consistent a model’s performance is across 10 randomly sampled, stratified subsets of the same sector, that is, its stability over similar data distributions. Specifically, we compute the mean and deviation of performance values relative to predictions made from the trend-prediction test set defined in Section 3.6.2.
Selected Model ($\alpha=0.5$): urf4_model_variant_n200_d13_c2_n50ppm
Target Performance Metrics (trend analysis results):
@@
  • Both metrics exhibit low variance (standard deviation < 10% of mean)
@@
Results: Evaluation across 10 random subsets ($\approx$40,000 total light curves) yields Table 2
Metric Value Target Deviation
Mean TOI Recall AUC 0.547 0.561 -0.014
Mean TOI Importance AUC 0.143 0.120 +0.023
Mean Binary TOI Recall 0.321 - -
Mean Anomaly Rate 0.279 - -
@@
Mean Importance (Top 20%) Average importance of TOIs within top 20% of scores Shows strength of early-ranked candidates - critical for targeted follow-up
Binary TOI Recall Fraction of TOIs flagged as anomalies (score > 0) Direct measure of detection breadth for candidate mining
Anomaly Rate Fraction of all light curves flagged as anomalous Reflects follow-up cost - high rates imply more manual vetting
Percentile Range (Top N%) Anomaly score percentile range containing top N% of TOIs by Importance Score Measures how early high-value objects appear - useful for threshold setting
@@
Interpretation: The evaluation demonstrates strong model generalization: TOI Recall AUC deviates only 0.014 below target, indicating consistent TOI prioritization across the anomaly score distribution. TOI Importance AUC exceeds the target by 0.023, suggesting the model performs even better than expected at detecting astrophysically significant targets. The binary recall of 32.1% and anomaly rate of 27.9% indicate closely related TOI Recall and Anomaly Rate as predicted by Figure 4. The value of these metrics will be improved by predicting over a finer mesh of synthetic design features and using a more comprehensive model scoring metric as presented in Section 4.1.
Verdict: This model satisfies our generalization criteria (that it will perform consistently on similar data), with both metrics falling within the acceptable ±0.03-0.05 deviation range. The empirical results confirm that the $\alpha$ = 0.5 design successfully balances recall and importance while maintaining stable performance across similar (in terms of ratio of TOIs to total curves) randomly stratified sampled data subsets from sector 2.
@@ 3.8.2 Hypothesis Validation: Alpha Variants
Goal: Assess whether changing $\alpha$ systematically alters model behavior across multiple evaluation dimensions - not just recall versus importance, but also prioritization sharpness, anomaly rate, and TOI distribution in score space. Model selection by this $\alpha$ metric is subject to a light supervisory step as referred to in Appendix A.5.
@@
URF-4 $\alpha$ variants are evaluated across six complementary metrics, each capturing different aspects of anomaly detection performance and scientific utility as in Table 3
3.8.3 Results: Alpha Variant Performance Comparison
Metric $\alpha$ = 0.3 $\alpha$ = 0.5 $\alpha$ = 0.9 Best Performer
TOI Recall AUC 0.521 0.547 0.557 $\alpha$ = 0.9
TOI Importance AUC 0.141 0.143 0.139 $\alpha$ = 0.5
Mean TOI Importance (Top 20%) 0.393 0.422 0.359 $\alpha$ = 0.5
Binary TOI Recall 0.493 0.321 0.682 $\alpha$ = 0.9
Anomaly Rate 0.447 0.279 0.558 $\alpha$ = 0.5
Metric S1 Best S2 Best S3 Best S4 Best S5 Best Consistency
3.8.4 Intepretation of Results of Table 4
  1. 1.
    $\alpha$-parameter control works as designed: Higher $\alpha$ values consistently increase recall-based metrics while lower $\alpha$ values enhance importance-based prioritization
  2. 2.
    Behavioral differences are robust: Performance patterns persist across all 10 evaluation subsets
  3. 3.
    Trade-offs are quantifiable: The six-metric framework captures the nuanced relationships between completeness, precision, and scientific value
@@ 3.8.5 Cross-Sector Generalization of Alpha Variants
Having established that $\alpha$ variants exhibit distinct and stable behavior within Sector 2, we next evaluate cross-sector generalization to test whether these behavioral patterns persist across different TESS observational periods. This analysis addresses a critical question for practical deployment: do models selected and characterized on one sector maintain their performance characteristics when applied to entirely different stellar populations and observing conditions?
@@
The means of each of the six metrics have been presented in Figure 9. The best performing alpha variant across the metrics have been presented in Table 5.
3.8.6 Results: Cross-Sector Generalization of Alpha Variants
Operational Implications: The predictable cross-sector behavior enables confident deployment of $\alpha$ variants to new TESS sectors without retraining, supporting mission-critical survey strategy decisions based on scientific priorities (set using a custom combined scoring system using performance metrics that represent that specific goals).
@@ 3.9 Benchmarking URF Models
To evaluate the impact of controllable synthetic design, we benchmark the performance of our URF-4 $\alpha$-variants against previously established baselines: URF-1 (uniform noise), URF-2 (fixed hyperparameters from MG23), and URF-3 (TOI-based synthetic set). This benchmarking provides a comprehensive view of the performance landscape for URF-based anomaly detection and situates CLARA’s synthetic design approach within it.
Limitations of Prior URF Variants:
  • URF-1: Overly conservative, achieving only 11% TOI recall. This model prioritizes purity but fails to identify most real transits.
  • URF-2: Based on hyperparameter settings from prior work. It demonstrates reasonable recall but lacks both flexibility and interpretability.
  • URF-3: Overly aggressive, labeling nearly 96% of light curves as anomalous. While it captures nearly all TOIs, it does so at the cost of overwhelming noise.
Advantage of URF-4 Synthetic Design: URF-4 circumvents these trade-offs by using physically motivated synthetic injections to enable $\alpha$-tunable control. Each $\alpha$ value corresponds to a scientific objective:
  • Alpha 0.3: Prioritizes importance-weighted metrics, ideal for high-confidence discovery pipelines.
  • Alpha 0.5: Provides an optimal balance—yielding 33% recall with only 27% anomaly rate..
  • Alpha 0.9: Maximizes completeness with high recall, suitable when follow-up resources are abundant.
Figure 10 shows a radar plot for Sector 1 comparing URF variants across normalized metrics including TOI recall, anomaly rate, importance-weighted metrics, and variability in top-importance scores. Ideal models minimize top_10p_by_importance_percentile_range_upper, strike a trade-off between toi_recall and anomaly_rate, and maximize mean_importance_top20p.
@@
Key Finding: Fixed hyperparameter models (URF-1, URF-2, URF-3) lock us into rigid, suboptimal behaviors. In contrast, our synthetic design enables targeted optimization for different scientific objectives - whether maximizing recall, minimizing false positives, or improving the astrophysical quality of discoveries. URF-4 $\alpha$-variants occupy the “Goldilocks zone” of anomaly detection: not too conservative, not too aggressive, but controllably optimized and generalizable over multiple sectors.
3.10 Conclusion (Part I)
The comprehensive evaluation of URF-4 $\alpha$ variants establishes the successful validation of our central hypothesis: synthetic set design enables systematic steering of anomaly detection models toward different scientific objectives. Cross-sector analysis across TESS Sectors 1-5 demonstrates robust behavioral consistency, with each $\alpha$ variant maintaining its designed priorities across diverse stellar populations and observing conditions.
Key findings include:
  • High-$\alpha$ models ($\alpha$ = 0.9) consistently maximize detection completeness, achieving best Binary TOI Recall across all sectors
  • Low-$\alpha$ models ($\alpha$ = 0.3) excel at importance-based prioritization, concentrating high-value targets within narrow anomaly score ranges
  • Balanced models ($\alpha$ = 0.5) provide optimal precision-recall trade-offs while maintaining the lowest false positive rates.
The predictable cross-sector generalization validates the transferability of synthetic training approaches and enables confident deployment to new TESS sectors without retraining.
Having established optimal URF architectures and validated behavioral steering mechanisms, we now proceed to comprehensive parameter space optimization to identify the single best-performing model configurations for $\alpha$=0.3 and $\alpha$=0.9. Through systematic evaluation across the following parameter grid:
@@ 3.10 Conclusion (Part I)
The next part of this paper validates the astrophysical significance of the optimized URF model through correlation analysis between anomaly scores and known stellar/planetary properties. Using improved Normalized Weighted Root Sum of Squares (N-WRSS) scoring and t-SNE clustering analysis, we demonstrate that high anomaly scores correspond to genuine astrophysical phenomena rather than instrumental artifacts. Morphological classification pipelines filter light curves based on SIMBAD catalog similarity, providing quantitative validation of URF detection capabilities.
4 Astrophysical Relevance and Morphological Clustering of URF Anomalies For Transit Detection
In this section, we shift focus from model selection and parameter optimization to the astrophysical validation and interpretability of anomalies detected by the CLARA framework. Having established that synthetic set design can systematically steer URF behavior, we now seek to answer a fundamental question: Do the anomalies identified by our optimized models correspond to genuine astrophysical phenomena, or are they dominated by noise and instrumental artifacts?
To address this first, we perform a feature-to-metric mapping over a finer mesh of synthetic set feature values as described in section 3.10, systematically varying synthetic light curve parameters (such as noise, and duration, as they were the most important feature according to the Random Forest Regressor) to pick 0.3 and 0.9 $\alpha$ variant urf-4 models as described in the previous section.
Second, we rigorously validate the astrophysical significance of detected anomalies. By correlating anomaly scores with known stellar and planetary properties, and employing techniques such as Normalized Weighted Root Sum of Squares (N-WRSS) scoring and t-SNE clustering (van der Maaten & Hinton, 2008), we map the distribution of high-scoring anomalies in both feature and astrophysical parameter space. Furthermore, we leverage external catalogs such as SIMBAD (Wenger et al., 2000) to perform morphological classification, quantitatively demonstrating that our models are sensitive to real transit-like events and astrophysically meaningful outliers.
@@ 4.1 A Finer Mapping of Features to Performance Metrics
These two parameters — noise (ppm) and transit duration — were selected for finer-grained grid variation based on feature importance analysis from earlier URF-4 performance modeling (Figures 67). Random Forest regressors revealed that noise most strongly influenced TOI Importance AUC, while duration was the dominant driver of anomaly rate and TOI Recall behavior. This empirical insight directly motivated our decision to fix all other variables and systematically vary only these two dimensions, enabling targeted exploration of their isolated effects on performance under the multi-metric $\alpha$-model selection criteria in Section 4.
The grid was intentionally densified in the ranges of duration $\sim$23–25.5 days and noise $\sim$150–300 ppm, where prior experiments indicated sharp performance transitions, enabling finer resolution of model behavior in these critical regions.
With fixed values of ncurves=300, duration(days A.3)=13 and cadence_minutes=2, we systematically varied the noise parameter in the range [25, 800]. To rank models based on varied values of noise (ppm), we used the following revised combined scoring metric:
$\text{Combined Score}=\alpha\cdot R+(1-\alpha)\cdot(1-I)-\beta\cdot A$ (6)
where a high value of $\beta$ penalizes anomaly rate $A$ more. The $\beta$ term was introduced due to significant differences between anomaly rate and recall at several values for noise as demonstrated in Figure 11. (1I)(1-I) to minimize $I$, which represents the upper limit of Top 10P Importance Range (range of anomaly score that contains the top 10 percent TOIs by importance score).
Noise (ppm) values 50e-6 and 225e-6 turned out most dominant in terms of importance. The 10Pimportance Upper Limit values for the most dominant noise values were 1.70 and 34.22 respectively. The TOI Recall values were 0.26 and 0.63 respectively. The Anomaly Rates were 0.29 and 0.58 respectively. We now set noise (ppm) at 50 and 225 and move on to varying values of duration, with more values in the 13 days to 27 days range. The following values were tested for duration (days A.3):
0.73, 1.65, 3.0, 3.58, 5.1, 6.0, 7.0, 8.3, 10.0, 12.0, 13.0, 15.67, 17.0, 20.0, 21.0, 21.5, 22.0, 22.5, 23.0, 23.25, 23.5, 23.75, 24.0, 24.25, 24.5, 24.75, 25.0, 25.25, 25.5, 26.0, 27.0
The range from 20 days to 26 days showed predictable variability for both noise values. (Noise (ppm) = 100 was also tested for the three duration values for $\alpha$=.3, $\alpha$=.5, and $\alpha$=.9 models) as shown in Figure 13 and 14 (figure 13 scaled up to 20 days to 27 days).
@@
$S_{\text{combined}}(\alpha)$ $=w_{\text{recall}}\cdot R_{\text{norm}}+w_{\text{anomaly}}^{-1}\cdot A_{\text{inv}}+w_{\delta}\cdot\Delta_{\text{norm}}$
$\quad+w_{\text{importance}}\cdot I_{\text{norm}}+w_{\text{range}}\cdot P_{\text{inv}}$ (7)

### v3

@@ 2.5 Controlled Parameter Space Exploration
URF-4’s framework enables systematic exploration of four key synthetic parameters (and more not explored ones):
@@
Each parameter combination from the cartesian product space of the parameter grid values is used to fine tune hyper-parameters (2.3) independent URF models against a fixed real light curve dataset, resulting in $3\times 2\times 2\times 3=36$ sub-variant models.
@@ 2.6.2 Controlled Subset for Model Comparison
  • Trend Analysis Step: The performance of the 36 models on this small subset helps us analyze trends in their behavior in relation to the parameters of synthetic set generation and quantify those trends using the performance metrics in 2.6.4 and then test the generalization of those trends on independent sectors using cross sector evaluation of the models with more (and better, in terms of representation of quantification of model performance) performance metrics described in Table 1
@@ 2.6.3 TOI Importance Weighting
where SNR represents the signal-to-noise ratio of the planet detection, Depth (ppm) is the transit depth in parts per million, and Duration (days) is the transit duration in days. This heuristic emphasizes TOIs that are deeper, longer-duration, and higher signal-to-noise, making them both more detectable and more reliable for anomaly detection evaluation. For fair comparison across sectors, TOI importance scores are normalized to the range $[0,1]$.
@@ 2.6.4 Performance Metrics
We emphasize that TOIs are never used in training or hyperparameter tuning. They serve solely as an independent performance metric. Cross-sector tests (Section 2.7.3) provide additional validation of model stability, while clustering and cosine similarity analyses (Section 2.9) are presented only for contextual interpretation, not model selection. This ensures that our evaluation avoids circular validation and remains consistent with the unsupervised framework.
@@ 2.7 Combined Metric Analysis
$\text{S}_{\text{combined}}$ $=\alpha\cdot\text{Recall AUC}_{\text{TOI}}$
$\quad+(1-\alpha)\cdot\text{Importance AUC}_{\text{TOI}}$ (3)
The weighting parameter $\alpha$ enables systematic control over the scientific trade-off between casting a wider net (high $\alpha$, emphasizing comprehensive TOI recall) versus casting a more accurate net (low $\alpha$, prioritizing astrophysically significant (higher values of importance metrics) targets).
@@
    Convexity: Since $\alpha\in[0,1]$ $S_{\text{combined}}$ lies within the convex hull of $R$ and II. This guarantees that the combined score is always bounded between the two metrics and cannot extrapolate beyond them i.e. $S_{\text{combined}}\in[\min(R,I),\max(R,I)]$.
@@
In Part II of this paper, we introduce a more comprehensive version of this combined score, adjusted for a total of six performance metrics, that satisfy the same properties as listed above.
@@
To validate the stability and generalization of our balanced URF-4 configuration for $\alpha$=0.5, we evaluate the balanced model across multiple random subsets of unseen data compiled with stratified sampling from the same sector (sector 2). In this context, by “generalization” we refer to how consistent a model’s performance is across 10 randomly sampled, stratified subsets of the same sector, that is, its stability over similar data distributions. Specifically, we compute the mean and deviation of performance metrics relative to predictions made from the trend-prediction test set defined in Section 2.6.2.
Selected Model ($\alpha=0.5$): urf4_n200_d13_c2_n50ppm
Target Performance Metrics (predicted from trend analysis results):
@@
  • Both metrics exhibit low variance from predicted values(deviation ¡ 10% of mean)
@@
Results of this generalization has been explored in Section 3.2.
@@
Mean Importance (Top 20%) Average importance of TOIs within top 20% of scores Shows strength of early-ranked candidates — critical for targeted follow-up
Binary TOI Recall Fraction of TOIs flagged as anomalies (score >> 0) Direct measure of detection breadth for candidate mining
Anomaly Rate Fraction of all light curves flagged as anomalous Reflects follow-up cost — high rates imply more manual vetting
Percentile Range (Top N%) Anomaly score percentile range containing top N% of TOIs by Importance Score Measures how early high-value objects appear — useful for threshold setting
@@ 2.7.2 Hypothesis Validation: Alpha Variants
Goal: Assess whether changing $\alpha$ systematically alters model behavior across multiple evaluation dimensions - not just recall versus importance, but also prioritization sharpness, anomaly rate, and TOI distribution in score space. We then extend this across five separate sectors to test generalization on independent populations. Model selection by this $\alpha$ metric is subject to a light supervisory step as referred to in Appendix B.5.
@@
URF-4 $\alpha$ variants are evaluated across six complementary metrics, each capturing different aspects of anomaly detection performance and scientific utility as described in Table 1. Results of this validation test for alpha variants working as per design have been discussed in Section 3.2.2.
@@ 2.7.3 Cross-Sector Generalization of Alpha Variants
We next evaluate cross-sector generalization to test whether these behavioral patterns (Section 3.2.2) persist across different TESS observational periods. This analysis addresses a critical question for practical deployment: do models selected and characterized on one sector maintain their performance characteristics when applied to entirely different stellar populations and observing conditions?
@@
The means of each of the six metrics have been presented in Figure 17. The best performing alpha variant across the metrics have been presented in Table 4. Results have been discussed in Section 3.2.3.
@@ 2.8 Benchmarking URF Models
To evaluate the impact of controllable synthetic design, we benchmark the performance of our URF-4 $\alpha$-variants against previously established baselines: URF-1 (uniform noise), URF-2 (fixed hyperparameters from MG23), and URF-3 (TOI-based synthetic set). Figure 2 shows a radar plot for Sector 1 comparing URF variants across normalized metrics including TOI recall, anomaly rate, importance-weighted metrics, and variability in top-importance scores. Ideal models minimize top_10p_by_importance_percentile_range_upper, strike a trade-off between toi_recall and anomaly_rate, and maximize mean_importance_top20p. As evident by the radar plot, URF-4 occupies the goldilocks zone compared to the performance of URF 1-3.
@@
2.9 Part 2 - Astrophysical Relevance and Morphological Clustering of URF Anomalies For Transit Detection
Having established optimal URF architectures and validated behavioral steering mechanisms, in this section, we shift focus from model selection and parameter optimization to the astrophysical validation and interpretability of anomalies detected by the CLARA framework. We now seek to answer a fundamental question: Do the anomalies identified by our optimized models correspond to genuine astrophysical phenomena, or are they dominated by noise and instrumental artifacts?
To address this first, we perform a feature-to-metric mapping over a finer mesh of synthetic set feature values as follows, systematically varying synthetic light curve parameters (such as noise, and duration, as they were the most important feature according to the Random Forest Regressor) to pick 0.3 and 0.9 $\alpha$ variant urf-4 models as described in the previous section.
@@ 2.9 Part 2 - Astrophysical Relevance and Morphological Clustering of URF Anomalies For Transit Detection
Second, we rigorously validate the astrophysical significance of detected anomalies. By correlating anomaly scores with known stellar and planetary properties, and employing techniques such as Normalized Weighted Root Sum of Squares (N-WRSS) scoring and t-SNE clustering (L. van der Maaten & G. Hinton, 2008), we map the distribution of high-scoring anomalies in both feature and astrophysical parameter space. Furthermore, we leverage external catalogs such as SIMBAD (M. Wenger et al., 2000b) to perform morphological classification, quantitatively demonstrating that our models are sensitive to real transit-like events and astrophysically meaningful outliers.
@@ 2.10 A Finer Mapping of Features to Performance Metrics
These two parameters — noise (ppm) and transit duration — were selected for finer-grained parameter grid variation tests (over the values mentioned in Section 2.9, based on feature importance analysis using the random forest regressor from earlier URF-4 performance modeling (Figures 56). The grid was intentionally densified in the ranges of duration $\sim$23–25.5 days and noise $\sim$150–300 ppm, where prior experiments indicated sharp performance transitions, enabling finer resolution of model behavior in these critical regions.
Noise (ppm) values 50e-6 and 225e-6 turned out most dominant in terms of importance. The 10Pimportance Upper Limit values for the most dominant noise values were 1.70 and 34.22 respectively. Fixing these values for noise, we varied duration for which the range from 20 days to 26 days showed predictable variability for both noise values. (Noise (ppm) = 100 was also tested for the three duration values for $\alpha$=.3, $\alpha$=.5, and $\alpha$=.9 models) as shown in Figure 22 and 3 (figure 22 scaled up to 20 days to 27 days).
@@
$S_{\text{combined}}(\alpha)$ $=w_{\text{recall}}\cdot R_{\text{norm}}+w_{\text{anomaly}}^{-1}\cdot A_{\text{inv}}+w_{\delta}$
$\quad\cdot\Delta_{\text{norm}}+w_{\text{importance}}\cdot I_{\text{norm}}+w_{\text{range}}\cdot P_{\text{inv}}$ (4)